OpenStudy (anonymous):

What is the equation of the line that passes through ( 1,5) and (4,11)?

OpenStudy (anonymous):

do you know what the equation for a line is?

OpenStudy (anonymous):

np

OpenStudy (anonymous):

What is the equation of the line that passes through ( 1,5) and (4,11)? Select one: a. y = 6 b. y = x + 2 c. y = 2x + 3 d. y = 3x + 4

OpenStudy (anonymous):

ok so what is the basic equation for any line

OpenStudy (anonymous):

y= bx+m

OpenStudy (anonymous):

or is it the other way around y= mx +b

OpenStudy (anonymous):

that is exactly right so do you know how what m is and how to calculate it

OpenStudy (anonymous):

is it the slope ?? and no

OpenStudy (anonymous):

ok so you are right it is the slope, so you calculate it with the following formulae: \[m=\frac{ (y2-y1) }{ (x2-x1) }\] I this may look complex but all but (x,y) is simply a coordinate so x1 is the x value of the first coordinate y 1 is the y value of the first coordinate

OpenStudy (ness9630):

First find the slope of the line: \[\LARGE \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] with \[\LARGE (x_{1}, y_{1});(x_{2}, y_{2})\]

OpenStudy (ness9630):

Once you fond the slope, use the equation: \[\LARGE y-y_{1}=m(x-x_{1})\] Where m=slope

OpenStudy (ness9630):

Then after using the point-slope formula get it into slope-intercept form: \[\LARGE y=mx+b\]

OpenStudy (anonymous):

sooo from the answr choices earlier the answer is A

OpenStudy (anonymous):

What is the equation of the line that passes through ( 1,5) and (4,11)? Select one: a. y = 6 b. y = x + 2 c. y = 2x + 3 d. y = 3x + 4

OpenStudy (ness9630):

We're not here to give you answers, we are here to guide you towards it

OpenStudy (anonymous):

NUT I DONT KNOW IT

OpenStudy (anonymous):

BUT